OpenStudy (anonymous):

How do I find the derivative of. 180 * (.5)^x

OpenStudy (anonymous):

Rewrite it as\[f(x)=(180 ) (.5)^x=180e^{xln(.5)}\]Let u=x ln(.5) and use the chain rule.

OpenStudy (anonymous):

So\[f'(x)=180 \ln(.5) e^{x \ln(.5)}=180 \ln(.5) (.5)^x\]

OpenStudy (anonymous):

Hope that was helpful.

OpenStudy (anonymous):

Thank you. I thought if I could do it with the rule of \[b^x = \ln(b)*b^x\] but the 180 threw me off.